OpenStudy (anonymous):

Trigonometric expression as an algebraic expression. I am having trouble sin(arctan u)

OpenStudy (dumbcow):

Let \[\theta = \tan^{-1} u\] then \[\tan \theta = \frac{u}{1}\] Draw a triangle  \[\sin \theta = ?\]

OpenStudy (anonymous):